OpenStudy (anonymous):

(x3 - 18x2 +95x - 126) ÷ (x -9)

OpenStudy (anonymous):

Long division - do u know how to do it?

OpenStudy (anonymous):

\[ x^3-18 x^2+95 x-126=(x-9) (x-7) (x-2) \]

OpenStudy (anonymous):

no

OpenStudy (anonymous):

okay, here's a site to show you: http://www.purplemath.com/modules/polydiv2.htm i would draw it out but i'm on a laptop right now so my hannswriting is barely legible.

OpenStudy (anonymous):

*handwriting

OpenStudy (anonymous):

\[ \frac{x^3-18 x^2+95 x-126}{x-9}=\frac{(x-9) (x-7) (x-2)}{x-9}=(x-7) (x-2) \] but, I suggest that you do it with using long division.

OpenStudy (anonymous):

is it x^2 + 9x + 14 or x^2 - 9x + 14

OpenStudy (anonymous):

yes, because that factors to (x-7)(x-2), good job!

OpenStudy (anonymous):

wait which one

OpenStudy (anonymous):

the second one

OpenStudy (anonymous):

Here how you factor the numerator: \[ x^3 - 18 x^2 + 95 x - 126=\\ x( x^2 - 18 x + 95) -126=\\ x( x^2 -18x + 81 -81 +95) -126=\\ x( x^2-9) + 14 x -126 =\\ x( x-9)^2 + 14 (x -9) =\\ (x-9)(x^2 -9x +14)= \\ (x-9)(x-7)(x-2) \]

OpenStudy (anonymous):

@eliassaab nice method, i'll try it sometime (: